# A Survey on Deep Learning based Time Series Analysis with Frequency Transformation

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series data is amongst the most ubiquitous data types, and has penetrated nearly every corner of our daily life (FDama_forecast_2021, ), e.g., user-item interaction series in e-commerce and stock price series over time in finance. In recent years, time series analysis has attracted rapidly increasing attention from academia and industry, particularly in areas such as time series forecasting (Benidis_forecasting_2022, ), anomaly detection (Zahra_anomaly_2022, ), and classification (Ismail_classification_2019, ). Time series analysis has played a critical role in a wide variety of real-world applications to address significant challenges around us long-lastingly, such as traffic monitoring (Bai2020nips, ), financial analysis (FengHWLLC19, ), and COVID-19 prediction (tampsgcnets2022, ). However, time series analysis is extremely challenging due to the intricate inter-series correlations and intra-series dependencies.
@@
Previous time series models based on deep learning have been devoted to modeling complex intra- and inter-series dependencies in the time domain to enhance downstream tasks. Representative sequential models such as recurrent neural networks (RNNs) (LSTNet_2018, ; LSTMNAD_2018, ), temporal convolutional networks (TCNs) (TCN_2018, ), and attention networks (autoformer_2021, ) are utilized to capture intra-series dependencies, while convolutional networks such as convolutional neural networks (CNNs) (DCRNN_2018, ) and graph neural networks (GNNs) (Chen_2022, ) are preferred to attend to inter-series correlations. Although achieving good results, those networks have inherent drawbacks of time-domain modeling, limiting their capabilities in capturing critical patterns for time series analysis. For example, GNNs are constructed based on variable-wise connections as illustrated in Fig. 1(a), and the sequential models (i.e., Transformer, RNN, and TCN) are based on timestamp-wise connections as shown in Fig. 1(b), (c), and (d), respectively. These modelings consider point-wise (e.g., variable/timestamp-wise) connections and fail to attend to whole or sub time series. Therefore, they are usually incapable of modeling common but complex global patterns, such as periodic patterns of seasonality, in time series (ATFN_2022, ; CoST_2022, ). These inherent drawbacks inspire researchers to address the intricate inter-series correlations and intra-series dependencies of time series from a different perspective.
@@
Recently, deep learning methods leveraging frequency transformation (FT) (roberts1987digital, ), e.g., Discrete Fourier Transform (DFT) (winograd1976computing, ), Discrete Cosine Transform (DCT) (ahmed1974discrete, ), and Discrete Wavelet Transform (DWT) (shensa1992discrete, ), have gained a surge of interest within the machine learning community (XuCVPR2020, ; ChiJM20, ; John2021, ; ZhouYZW22, ). These neural models incorporating frequency transformation have demonstrated an efficient learning paradigm in time series analysis and achieved state-of-the-art performance in terms of both efficiency and effectiveness (autoformer_2021, ; fedformer_2022, ; TFAD_2022, ). This can be attributed to the distinctive advantages of FT (see Section 6.1) that the frequency spectrums generated by FT contain abundant vital patterns, e.g., seasonal trends, and provide a global view of the characteristics of time series. In addition, FT facilitates obtaining multi-scale representations and multi-frequency components of time series for capturing informative representations and patterns. This motivates us to systematically summarize and analyze the advantages of FT to instruct researchers in this area and to deliver a comprehensive survey on the emerging area, i.e., deep learning based time series analysis with FT, thereby enlightening the time series community. While the literature includes various studies that discuss time series analysis from different perspectives (Fakhrazari_2017, ; Benidis_forecasting_2022, ; FawazFWIM19, ; ChenPZS21, ; ChenKS22, ; 0001EL21, ; liang_kdd2024, ; zhang_pami2024, ; Wang_arxiv_2024, ), there remains a lack of comprehensive summaries on the topic of time series analysis with FT. Moreover, the reasons why FT can enhance the time series analysis have not yet been summarized, and its limitations have not been thoroughly analyzed. These gaps have hindered the theoretical development and practical applications of time series analysis with FT.
@@
The subsequent sections of this paper are structured as follows: Section 2 initially presents the fundamental concepts of time series analysis and frequency transformation. Following that, Section 3 summarizes existing FT-equipped models in terms of approaches incorporating FT to enhance the accuracy or efficiency of time series analysis. In Section 4, we delve into the practical implementations of these models and examine the types of neural works utilized in conjunction with FT. Subsequently, Section 5 categorizes representative frequency-based methods based on common time series tasks, including forecasting, anomaly detection, classification, and imputation. In Section 6, we discuss the advantages and limitations of the frequency domain. Finally, we enlighten new avenues of future directions for time series analysis in Section 7.
@@ 2.1.1. Forecasting
Time series forecasting is the task of extrapolating time series into the future (Benidis_forecasting_2022, ). For a given time series $\mathbf{X}=[{X}_{1},{X}_{2},\cdots,{X}_{T}]\in\mathbb{R}^{N\times T}$ with $N$ series and $T$ timestamps, where ${X}_{t}\in\mathbb{R}^{N}$ denotes the multi-variate values of $N$ distinct series at timestamp $t$. We consider a time series lookback window of length-$L$ at timestamp $t$, namely $\mathbf{X}_{t}=[{X}_{t-L+1},{X}_{t-L+2},\cdots,{X}_{t}]\in\mathbb{R}^{N\times L}$; also, we consider a horizon window of length-$\tau$ at timestamp $t$ as the prediction target, denoted as $\mathbf{Y}_{t}=[{X}_{t+1},{X}_{t+2},\cdots,{X}_{t+\tau}]\in\mathbb{R}^{N\times\tau}$. Then the time series forecasting task is to use historical observations $\mathbf{X}_{t}$ to predict future values $\hat{\mathbf{Y}}_{t}$ and the typical forecasting model $f_{\theta}$ parameterized by $\theta$ is to produce forecasting results by $\hat{\mathbf{Y}}_{t}=f_{\theta}(\mathbf{X}_{t})$.
@@ 2.1.2. Classification
Time series classification seeks to assign labels to each series of a dataset (FawazFWIM19, ). Generally, for a time series dataset $D=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\cdots,(X_{N},Y_{N})\}$ where $X_{i}\in\mathbb{R}^{T}$ is a time series with $T$ timestamps and $Y_{i}$ is its corresponding one-hot vector label. For the dataset $D$ containing $K$ classes, $Y_{i}$ is a vector of length $K$ where each element $j\in[1,K]$ equals 1 if the class of $X_{i}$ is $j$ and 0 otherwise. Then the classification task is to train a classifier $f_{\theta}$ parameterized by $\theta$ on the dataset $D$ to map from the space of possible inputs to a probability distribution over the class variable labels, formulated as $Y_{i}=f_{\theta}(X_{i})$.
@@ 2.1.3. Anomaly Detection
Time series anomaly detection seeks to find abnormal subsequences in a series (ChenPZS21, ). The goal is to develop algorithms or models that can effectively distinguish between normal and anomalous behavior, thereby providing early detection and alerting for unusual events or behaviors in the time series data. Given a time series $X=[x_{1},x_{2},\cdots,x_{T}]$ with $T$ timestamps where $x_{i}$ represents the data point at time index $i$, the anomaly detection is to identify a subset of data points $X_{s}\subseteq X$ that represents the anomalous or abnormal instances.
@@ 2.2.1. Discrete Fourier Transform
(2) $\mathcal{X}=\operatorname{Re}+j\operatorname{Im}$
@@
(3) $A=\sqrt{\operatorname{Re}^{2}+\operatorname{Im}^{2}}$
(4) $\theta=\arctan(\frac{\operatorname{Im}}{\operatorname{Re}})$
@@ 2.2.2. Discrete Cosine Transform
Discrete Cosine Transform (DCT) has emerged as the de-facto image transformation in most visual systems. The most common 1-D DCT $C(k)$ of a data sequence $x[n]$ is defined as
@@
where $k=0,1,...,N-1$, and $\alpha(k)$ is defined as
@@
DCT only retains the real parts of DFT and is roughly equivalent to DFT that has twice its length. It often performs on real data with even symmetry or in some variants where the input or output data are shifted by half a sample.
@@ 2.2.3. Discrete Wavelet Transform
(7) $\operatorname{DWT}(a,b)=a_{0}^{-m/2}\int_{-\infty}^{\infty}x(t)\Psi(a_{0}^{-m}t-kb_{0})(t)\mathrm{d}t$
In contrast to $\operatorname{DFT}$ and $\operatorname{DCT}$, a wavelet transform has the ability to identify the locations containing observed frequency content, while the DFT and DCT can only extract pure frequencies from the signal. Hence, $\operatorname{DWT}$ can perform time-frequency analysis. In addition, $\operatorname{DWT}$ can obtain different resolution representations (Mallat_1989, ) by changing the scaling and translation factors. In Table 1, we compare the three frequency transformation methods, including their pros and cons.
@@ 2.2.4. Convolution Theorem
The convolution theorem (1990Continuous, ) states that the Fourier transform of a circular convolution of two signals equals the point-wise product of their Fourier transforms. Given a signal $x[n]$ and a filter $h[n]$, the convolution theorem can be defined as follows:
@@
According to the Convolution Theorem, the point-wise product of frequency spectrums of two sequences is equivalent to their circular convolution in the time domain, where the product with a larger receptive field of the whole sequences better captures the overall characteristics (e.g., periodicity) and requires less computation cost (AlaaCS21, ).
@@ 3.1. Feature Engineering
Compared to the time domain, the frequency domain can provide vital information for time series, such as periodic information. Prior models take advantage of frequency domain information for periodic analysis and use it as an important complement to the time domain information. ATFN (ATFN_2022, ) proposes a frequency-domain block to capture dynamic and complicated periodic patterns of time series data, and integrates deep learning networks with frequency patterns. TFAD (TFAD_2022, ) utilizes a frequency domain analysis branch to detect complex pattern anomalies, e.g., periodic anomalies. CoST (CoST_2022, ) learns the trend representations in the time domain, whereas the seasonal representations are learned by a Fourier layer in the frequency domain. FreDo (fredo2022, ) is a frequency domain-based neural network model that is built on top of the baseline model to enhance its performance. ETSformer (etsformer2022, ) utilized DFT to design a frequency attention mechanism to replace the self-attention mechanism to identify seasonal patterns. DeepTime (WooLSKH23, ) leverages a novel concatenated Fourier features module to efficiently learn high-frequency patterns in time series.
@@
One big challenge for time series analysis is that there are intricate entangled temporal dynamics among time series data. To address this challenge, some methods try to solve it in terms of the frequency domain. They disentangle temporal patterns by decomposing time series data into different frequency components. SFM (SFM_2017, ) separates the memory states of RNN into different frequency states such that they can explicitly learn the dependencies of both the low- and high-frequency patterns. (ZhangAQ_kdd_2017, ) explicitly decomposes trading patterns into various frequency components and each component models a particular frequency of latent trading pattern underlying the fluctuation of stock price. Recently, wavelet-based models have shown competitive performances since wavelet transform can retain both time and frequency information and obtain multi-resolution representations. (WangWLW_kdd_2018, ) proposes a wavelet-based neural network structure for building frequency-aware deep learning models for time series analysis. (Wen_sigmod_2021, ) applies maximal overlap discrete wavelet transform to decouple time series into multiple levels of wavelet coefficients and then detect single periodicity at each level. (WangYJ023, ) devises a novel data-dependent wavelet attention mechanism for dynamic frequency analysis of non-stationary time series analysis. (YangLWZPW23, ) proposes an end-to-end graph enhanced Wavelet learning framework for long sequence forecasting which utilizes DWT to represent MTS in the wavelet domain.
@@
Existing time domain methods construct their models based on point-wise connections (see Fig. 1), which prevent them from capturing series-level patterns, such as overall characteristics of time series. By leveraging the global view property of the frequency domain, some works utilize frequency information to attend to series-level patterns. FEDformer (fedformer_2022, ) combines Fourier analysis with the Transformer which helps the Transformer better capture the global properties of time series. TFAD (TFAD_2022, ) integrates the frequency domain analysis branch with the time domain analysis branch and detects seasonality anomalies in the frequency domain. Besides, some works introduce frequency domain analysis to improve neural networks in order to address their inherent drawbacks. Vanilla convolutions in modern deep networks are known to operate locally, which causes low efficacy in connecting two distant locations in the network. To mitigate the locality limitation of convolutions, SRL (Chi_2019, ) converts data into the frequency domain and proposes spectral residual learning for achieving a fully global receptive field, and FFC (ChiJM20, ) harnesses the Fourier spectral theory and designs an operation unit to leverage frequency information for enlarging the receptive field of vanilla convolutions.

### v6

@@ Abstract.
*Corresponding Author
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736571isbn: 979-8-4007-1454-2/2025/08ccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series data is amongst the most ubiquitous data types, and has penetrated nearly every corner of our daily life (Dama and Sinoquet, 2021), e.g., user-item interaction series in e-commerce and stock price series over time in finance. In recent years, time series analysis has attracted rapidly increasing attention from academia and industry, particularly in areas such as time series forecasting (Benidis et al., 2022), anomaly detection (Darban et al., 2022), and classification (Fawaz et al., 2019). Time series analysis has played a critical role in a wide variety of real-world applications to address significant challenges around us long-lastingly, such as traffic monitoring (Bai et al., 2020; Yi et al., 2024b), financial analysis (Feng et al., 2019; Fei et al., 2025), and COVID-19 prediction (Chen et al., 2022c; He et al., 2025). However, time series analysis is extremely challenging due to the intricate inter-series correlations and intra-series dependencies.
@@
Previous time series models based on deep learning have been devoted to modeling complex intra- and inter-series dependencies in the time domain to enhance downstream tasks. Representative sequential models such as recurrent neural networks (RNNs) (Lai et al., 2018; Hundman et al., 2018), temporal convolutional networks (TCNs) (Bai et al., 2018), and attention networks (Wu et al., 2021) are utilized to capture intra-series dependencies, while convolutional networks such as convolutional neural networks (CNNs) (Li et al., 2018) and graph neural networks (GNNs) (Chen et al., 2022a) are preferred to attend to inter-series correlations. Although achieving good results, those networks have inherent drawbacks of time-domain modeling, limiting their capabilities in capturing critical patterns for time series analysis. For example, GNNs are constructed based on variable-wise connections as illustrated in Fig. 1(a), and the sequential models (i.e., Transformer, RNN, and TCN) are based on timestamp-wise connections as shown in Fig. 1(b), (c), and (d), respectively. These modelings consider point-wise (e.g., variable/timestamp-wise) connections and fail to attend to whole or sub time series. Therefore, they are usually incapable of modeling common but complex global patterns, such as periodic patterns of seasonality, in time series (Yang et al., 2022; Woo et al., 2022a). These inherent drawbacks inspire researchers to address the intricate inter-series correlations and intra-series dependencies of time series from a different perspective.
@@
Recently, deep learning methods leveraging frequency transformation (FT) (Roberts and Mullis, 1987), e.g., Discrete Fourier Transform (DFT) (Winograd, 1976), Discrete Cosine Transform (DCT) (Ahmed et al., 1974), and Discrete Wavelet Transform (DWT) (Shensa et al., 1992), have gained a surge of interest within the machine learning community (Xu et al., 2020; Chi et al., 2020; Guibas et al., 2022; Zhou et al., 2022c). These neural models incorporating frequency transformation have demonstrated an efficient learning paradigm in time series analysis and achieved state-of-the-art performance in terms of both efficiency and effectiveness (Wu et al., 2021; Zhou et al., 2022b; Zhang et al., 2022b). This can be attributed to the distinctive advantages of FT (see Section 6.1) that the frequency spectrums generated by FT contain abundant vital patterns, e.g., seasonal trends, and provide a global view of the characteristics of time series. In addition, FT facilitates obtaining multi-scale representations and multi-frequency components of time series for capturing informative representations and patterns. This motivates us to systematically summarize and analyze the advantages of FT to instruct researchers in this area and to deliver a comprehensive survey on the emerging area, i.e., deep learning based time series analysis with FT, thereby enlightening the time series community. While the literature includes various studies that discuss time series analysis from different perspectives (Fakhrazari and Vakilzadian, 2017; Benidis et al., 2022; Fawaz et al., 2019; Chen et al., 2021, 2022b; Schäfer et al., 2021; Liang et al., 2024; Zhang et al., 2024; Wang et al., 2024b), there remains a lack of comprehensive summaries on the topic of time series analysis with FT. Moreover, the reasons why FT can enhance the time series analysis have not yet been summarized, and its limitations have not been thoroughly analyzed. These gaps have hindered the theoretical development and practical applications of time series analysis with FT.
@@ 2.1.1. Forecasting
Time series forecasting is the task of predicting future data points based on past observations (Benidis et al., 2022). Consider a dataset with $N$ different time series measured over $T$ time steps, represented as $\mathbf{X}=[X_{1},X_{2},...,X_{T}]\in\mathbb{R}^{N\times T}$, where each $X_{t}\in\mathbb{R}^{N}$ captures the values of all series at time $t$. To make predictions at time $t$, we use a window of past $L$ observations, $\mathbf{X}_{t}=[X_{t-L+1},...,X_{t}]$, as input. The goal is to forecast the next $\tau$ values, denoted $\mathbf{Y}_{t}=[X_{t+1},...,X_{t+\tau}]$. A forecasting model, $f_{\theta}$, learns to map $\mathbf{X}_{t}$ to an estimate $\hat{\mathbf{Y}}_{t}=f_{\theta}(\mathbf{X}_{t})$.
@@ 2.1.2. Classification
Time series classification aims to assign a categorical label to each time series in a dataset (Fawaz et al., 2019). Formally, consider a dataset $D=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{N},Y_{N})\}$, where each $X_{i}\in\mathbb{R}^{T}$ represents a univariate time series with $T$ time steps, and $Y_{i}$ is its associated one-hot encoded label. Assuming $D$ contains $K$ distinct classes, each label $Y_{i}\in\mathbb{R}^{K}$ is a binary vector such that the $j$-th element is 1 if $X_{i}$ belongs to class $j$, and 0 otherwise. The goal is to learn a classifier $f_{\theta}$, parameterized by $\theta$, that maps input time series to class probability distributions: $Y_{i}=f_{\theta}(X_{i})$.
@@ 2.1.3. Anomaly Detection
Time series anomaly detection aims to identify abnormal patterns or subsequences within a temporal sequence (Chen et al., 2021). The objective is to design models or algorithms capable of distinguishing between normal and anomalous behaviors, enabling early detection and timely alerts for unusual events. Given a time series $X=[x_{1},x_{2},\ldots,x_{T}]$ with $T$ timestamps, where $x_{i}$ denotes the observation at time $i$, the task is to identify a subset $X_{s}\subseteq X$ corresponding to anomalous data points that deviate significantly from expected temporal patterns.
@@ 2.2.1. Discrete Fourier Transform
(2) $\mathcal{X}=\operatorname{Re}+j\operatorname{Im}.$
@@
(3) $A=\sqrt{\operatorname{Re}^{2}+\operatorname{Im}^{2}}.$
(4) $\theta=\arctan(\frac{\operatorname{Im}}{\operatorname{Re}}).$
@@ 2.2.2. Discrete Cosine Transform
Discrete Cosine Transform (DCT) has emerged as the de-facto image transformation in most visual systems. The most common 1-D DCT $C(k)$ of a data sequence $x[n]$ is defined as:
@@
where $k=0,1,...,N-1$, and $\alpha(k)$ is defined as:
@@
DCT only retains the real parts of DFT, and often performs on real data with even symmetry or in some variants where the input or output data are shifted by half a sample.
@@ 2.2.3. Discrete Wavelet Transform
(7) $\operatorname{DWT}(a,b)=a_{0}^{-m/2}\int_{-\infty}^{\infty}x(t)\Psi(a_{0}^{-m}t-kb_{0})(t)\mathrm{d}t.$
In contrast to $\operatorname{DFT}$ and $\operatorname{DCT}$, $\operatorname{DWT}$ has the ability to identify the locations containing observed frequency content, while the DFT and DCT can only extract pure frequencies from the signal. Hence, $\operatorname{DWT}$ can perform time-frequency analysis. In addition, $\operatorname{DWT}$ can obtain different resolution representations (Mallat, 1989) by changing the scaling and translation factors. In Table 1, we compare the three frequency transformation methods, including their pros and cons.
@@ 2.2.4. Convolution Theorem
The convolution theorem (Soliman and Srinath, 1990) states that the Fourier transform of a circular convolution of two signals equals the point-wise product of their Fourier transforms. Given a signal $x[n]$ and a filter $h[n]$, the convolution theorem can be defined as follows:
@@
According to the Convolution Theorem, the point-wise multiplication of the frequency spectra of two sequences corresponds to their circular convolution in the time domain. This operation, which inherently spans the entire sequence, enables more effective capture of global patterns such as periodicity, while also reducing computational cost (Alaa et al., 2021).
@@ 3.1. Feature Engineering
Compared to the time domain, the frequency domain can provide vital information for time series, such as periodic information. Prior models take advantage of frequency domain information for periodic analysis and use it as an important complement to the time domain information. ATFN (Yang et al., 2022) proposes a frequency-domain block to capture dynamic and complicated periodic patterns of time series data, and integrates deep learning networks with frequency patterns. TFAD (Zhang et al., 2022b) utilizes a frequency domain analysis branch to detect complex pattern anomalies, e.g., periodic anomalies. CoST (Woo et al., 2022a) learns the trend representations in the time domain, whereas the seasonal representations are learned by a Fourier layer in the frequency domain. FreDo (Sun and Boning, 2022) is a frequency domain-based neural network model that is built on top of the baseline model to enhance its performance. ETSformer (Woo et al., 2022b) utilized DFT to design a frequency attention mechanism to replace the self-attention mechanism to identify seasonal patterns. DeepTime (Woo et al., 2023) leverages a novel concatenated Fourier features module to efficiently learn high-frequency patterns in time series.
@@
One big challenge for time series analysis is that there are intricate entangled temporal dynamics among time series data. To address this challenge, some methods try to solve it in terms of the frequency domain. They disentangle temporal patterns by decomposing time series data into different frequency components. SFM (Hu and Qi, 2017) separates the memory states of RNN into different frequency states such that they can explicitly learn the dependencies of both the low- and high-frequency patterns. (Zhang et al., 2017) explicitly decomposes trading patterns into various frequency components and each component models a particular frequency of latent trading pattern underlying the fluctuation of stock price. Recently, wavelet-based models have shown competitive performances since wavelet transform can retain both time and frequency information and obtain multi-resolution representations. (Wang et al., 2018) proposes a wavelet-based neural network structure for building frequency-aware deep learning models for time series analysis. (Wen et al., 2021a) applies maximal overlap discrete wavelet transform to decouple time series into multiple levels of wavelet coefficients and then detect single periodicity at each level. (Wang et al., 2023) devises a novel data-dependent wavelet attention mechanism for dynamic frequency analysis of non-stationary time series analysis. (Yang et al., 2023) proposes an end-to-end graph enhanced Wavelet learning framework for long sequence forecasting which utilizes DWT to represent MTS in the wavelet domain.
@@
Existing time domain methods construct their models based on point-wise connections (see Fig. 1), which prevent them from capturing series-level patterns, such as overall characteristics of time series. By leveraging the global view property of the frequency domain, some works utilize frequency information to attend to series-level patterns. FEDformer (Zhou et al., 2022b) combines Fourier analysis with the Transformer which helps the Transformer better capture the global properties of time series. TFAD (Zhang et al., 2022b) integrates the frequency domain analysis branch with the time domain analysis branch and detects seasonality anomalies in the frequency domain. Besides, some works introduce frequency domain analysis to improve neural networks in order to address their inherent drawbacks. Vanilla convolutions in modern deep networks are known to operate locally, which causes low efficacy in connecting two distant locations in the network. To mitigate the locality limitation of convolutions, SRL (Chi et al., 2019) converts data into the frequency domain and proposes spectral residual learning for achieving a fully global receptive field, and FFC (Chi et al., 2020) harnesses the Fourier spectral theory and designs an operation unit to leverage frequency information for enlarging the receptive field of vanilla convolutions.